Fox Corporation has the following information for August: Beginning Work in Process Inventory (70% complete as to conversion) 6,000 units, started 24,000 units, Ending Work in Process Inventory (10% complete as to conversion) 8,500 units. Beginning WIP Inventory Costs: Material $23,400, Conversion $50,607. Current Period Costs: Material $31,500, Conversion $76,956. All material is added at the start of the process and all finished products are transferred out.  

Refer to Fox Corporation. Assume that FIFO process costing is used. What is the cost per equivalent unit for conversion? a)$3.44 b) $4.24 c) $5.71 d) $7.03
Answer #1

Equivalent units for conversion calculation-

Opening units are 70% complete with respect to conversion cost. Remaining 30% should be done in current period.

Opening units 6,000 X 30% = 1,800 units (A)

Units started = 24,000 units

Ending workin progress = 8,500 units

Units completed = 24,000 units - 8,500 units = 15,500 units (B)

Ending units are 10% complete as to conversion cost = 8,500 units X 10% = 850 units (C)

Equivalent units as to conversion cost = (A) + (B) + (C) = 1,800 units + 15,500 units + 850 units

= 18,150 units (D)

Conversion cost for the current period = $76,956 (E)

(E) / (D) = $76,956 / 18,150 units = $4.24

OPTION B